omework: Module 1 Homework 0 of 10 pts 2.9
Answer #1

class width :-According to the definition, the class width is the difference between two consecutive lower class limits in a frequency distribution.In the given table, the consecutive lower class limits are 5,7,9,11,13 so the class width is 7-5 =2.

class midpoints;:-

Class Midpoint: class midpoints are the values in the middle of the classes.These are computed by adding the lower class limit to the upper class limit and dividing the sum by 2.

Hence (5+6.9)/2=5.95 ,

(7+8.9)/2=7.95

similiarly we get 9.95, 11.95,and 13.95.

Class boundries:-The lower class boundary is found by subtracting 0.05 from the lower class limit and the upper class boundary is found by adding 0.05 to the upper class .

so, lower class boundries are 4.95,6.95,8.95,10.95,12.95

and upper class boundribou are 6.95,8.95,10.95,12.95,14.95